if v=3i - 4j and w= xi +9j, find all numbers x for which v+w=7
Answer #1

Ans: uz zinuj We xi +91 VtW e siruj txi +99 = (3 + x) 1+ 5j.. HIVI = Jazub 11 w+will - V (84x9+(5)? - Visi cx ??+ 3.3.x + 25-V6E+ (-8+216))*+6592 a scavot (52 arc4x6 ) +85 - VQ4+25 3/49 - at so , x=-3+285 is valid. - 18+(-8-ast))+ oga -E-206) ²+ (5)